Systems Staff Meeting Notes 9/20/99

Systems Staff Meeting Notes
September 20, 1999

Present: Rich Bennett, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. Network Coordinator position will remain posted until the end of this month.
  2. Water - Library Administration canceled their contract with Culligan. Systems will go back to buying our own water. If you are interested in chipping in on the water purchase, please let Debra know.
  3. Mail delivery problems to be discussed with Facilities at a meeting this afternoon.
  4. Storage - lack of storage space due to the multitude of equipment waiting for survey should be resolved today. Room 2 is full, paperwork is at a standstill - Rob Roberts says there should be no problem in getting it out.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Office 2000 - received this morning, is not checked in yet. Systems will be the first to upgrade. If we run across departments that are trying to upgrade from Office '97 to 2000, we will not assist them. Everyone should be on Office'98 at this point. If we receive any calls of this nature, they should be referred directly to Bill.
Liaisons
  1. Training Review session ended - will probably try to go with an in-house session next time. Reviews graded the program with an 'A' for effort, but negative for content. We may go with an outside company to bring in equipment or for use of an outside facility.
  2. There will probably be a class for the current liaisons on local networking issues (most likely presented by Systems). It will most likely coincide with the new liaison orientation to be held the first week October.
Networking
  1. IBM has not shown up yet. Two of the three control units in West are down following the post-Floyd startup. The floppies were suspect at first - now it seems more likely to be the disk drives. IBM is trying to find parts, but there's no guarantee with such antiquated equipment. The MSL controllers were flaky but eventually came up - another shutdown could kill them off as well.
  2. Other post-Floyd assessments - overall it went reasonably well, some departments did not comply - but some of those were departments with no active liaisons.
  3. SNA terminals are not Y2K compliant. Even the maintenance on the 3270 interface to LUIS is in question. LUIS may not even remain in place. Upgrading the necessary equipment would be costly. Need to be off of the SNA's by the end of the year.
  4. Mead Library just converted from the TN3270 this weekend. Iona was inquiring about returning the old equipment - all of the property should have been transferred to them. Journalism is almost completely off SNA's.
  5. Current Projects - network documentation needs to be moved to the top of the queue in priority. Needs to be completed within the next 2-3 days to facilitate the installation of the network switches. Other Systems staff may have to back Bobbie or Everett up in dealing with other computer problems while they work on the documentation.
Software Projects Status
  1. IR database - Ying is working on some new requirements - trying to include some functions of inventory by adding tables.
  2. Mellon2 - Winston created an MS Access table for David, but has not received feedback. This will be a data analysis project.
  3. OPS Tracker reports are finished & in use.
  4. Hilton has finished & installed the travel updates.
  5. Ying has been working on Configuration Management Duties, which is a parallel project to the network side of documentation. It acts as a virtual file cabinet for housing information about current versions, installation notes, problems, etc.
  6. Winston has been trying to decommission the SQL server on NT7 & move it to NT4. He still has 3 or 4 more applications to go. NT4 is currently housed in Martha's old office. NT3 will have to be briefly moved as well. There is talk of ultimately moving them to what is currently Grace's office.
DP Coordinator
  1. Compustat tape exchange still on hold pending CRSP Link CD purchase.
Calendar
  1. Bobbie out Wednesday, October 6.
  2. Suzy & Bill out October 20-November 1.
  3. Everett out October 22-November 1.


Back to the Systems Homepage

Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu